A car accelerates uniformly from rest and reaches a speed of 18.7 m/s in 6.3 s. The diameter of a tire is 39.4 cm.

Find the number of revolutions the tire makes during this motion, assuming no slip- ping.

Answer in units of rev.


1
Expert's answer
2022-01-19T06:01:12-0500

"s=\frac{(0+v)t}2=\frac {vt}2,"

"n=\frac s{\pi d}=\frac{vt}{2\pi d}=47.6."